# How to Complete Domain Tuning while Keeping General Ability in LLM: Adaptive Layer-wise and Element-wise Regularization

### v1

@@ Abstract
Large Language Models (LLMs) exhibit strong general-purpose language capabilities. However, fine-tuning these models on domain-specific tasks often leads to catastrophic forgetting, where the model overwrites or loses essential knowledge acquired during pretraining. This phenomenon significantly limits the broader applicability of LLMs. To address this challenge, we propose a novel approach to compute the element-wise importance of model parameters crucial for preserving general knowledge during fine-tuning. Our method utilizes a dual-objective optimization strategy: (1) regularization loss to retain the parameter crucial for general knowledge; (2) cross-entropy loss to adapt to domain-specific tasks. Additionally, we introduce layer-wise coefficients to account for the varying contributions of different layers, dynamically balancing the dual-objective optimization. Extensive experiments on scientific, medical, and physical tasks using GPT-J and LLaMA-3 demonstrate that our approach mitigates catastrophic forgetting while enhancing model adaptability. Compared to previous methods, our solution is approximately 20 times faster and requires only 10%–15% of the storage, highlighting the practical efficiency. The code will be released.
@@ 1 Introduction
Large Language Models (LLMs) are pretrained on massive and diverse datasets, equipping them with remarkable general-purpose language capabilities. This pretraining process allows LLMs to serve as versatile tools for a wide range of natural language processing tasks. However, in real-world applications, users often fine-tune these models for specific domains and tasks, such as adapting them to the medical or scientific domains. While fine-tuning could enhance the model task-specific performance, it also introduces a critical challenge: catastrophic forgetting. As shown in Figure 1, catastrophic forgetting refers to the phenomenon where a model, during the process of fine-tuning, loses or overwrites knowledge learned during pretraining. This occurs because the new domain-specific training data typically does not include the diverse and representative information present in the pretraining corpus. Consequently, gradients from the fine-tuning stage disproportionately update parameters critical to the general capabilities of LLMs, leading to a significant degradation in their performance on tasks outside the fine-tuned domain. This issue poses a severe limitation on the broader applicability of LLMs, as it undermines their versatility and reusability across domains.
@@
Addressing catastrophic forgetting is therefore a crucial requirement for maximizing the utility of LLMs. A successful solution needs to achieve a delicate balance: retaining the essential general-purpose knowledge when learning new domain-specific expertise. This balance is critical when fine-tuning LLMs for specialized tasks, as both domain adaptation and generalizability are necessary for practical applications. EWCLoRA (Xiang et al., 2024) focuses on the issue of catastrophic forgetting in LLM fine-tuning and uses the Fisher matrix to measure the importance of parameters for general capabilities. However, its computational cost is very high. For GPT-J-6B, calculating the Fisher matrix takes 22 hours on an A800 and requires 23GB of storage, and these requirements increase for larger LLMs. Besides, RsLoRA (Kalajdzievski, 2023) aims to stabilize learning by introducing a rank-stabilized scaling factor, but it does not effectively protect general capabilities as expected.
@@
To address the catastrophic forgetting, inspired by Synaptic Intelligence (SI) (Zenke et al., 2017), we propose a method to compute the element-wise importance of parameters within LLMs on general-purpose capabilities, preserving those parameters that are critical for general knowledge to minimize significant modifications to general knowledge. Specifically, we define domain $\nu$ as the general knowledge, representing the general-purpose capabilities of LLMs, and domain $\mu$ as the knowledge learned during fine-tuning for specific tasks. Our approach leverages a dual-objective optimization strategy that combines two loss functions: Regularization Loss, which reduces updates to parameters critical for domain $\nu$ to preserve general knowledge; and Cross-Entropy (CE) Loss, which encourages alignment of domain $\mu$ parameters to enhance domain-specific learning. Minimizing the regularization loss helps preserve general capabilities while minimizing the CE loss allows the model to fully adapt to the new domain.
@@
Besides, we propose a layer-wise coefficient to adjust the weight of regularization loss. In LLMs, different layers contribute differently to generalization ability and domain-specific ability. The impact of each layer on the learning process is not uniform; some layers capture high-level abstract features, while others focus on lower-level details. Traditional approaches often treat the importance of each layer as equal, which overlooks the varying degrees of influence that different layers have on the model learning and generalization ability. Thus, we propose layer-wise coefficients to dynamically adjust the balance between task learning and the retention of general knowledge in each layer, allowing some layers to prioritize task learning, while others preserve general knowledge. We use the L2 norm of the computed element-wise importance as a measure for each layer weight, as it captures the contribution of parameters of each layer to both objectives.
@@
  • We introduce a framework that first trains on general data and then fine-tunes on domain-specific data, effectively addressing catastrophic forgetting in large language models (LLMs).
@@
  • Our method achieves state-of-the-art performance across multiple datasets using mainstream backbone LLMs. It also significantly reduces computational time and storage requirements for parameter importance estimation compared to prior methods, demonstrating superior efficiency and scalability.
@@ 2 Related Work
2.1 Traditional Catastrophic Forgetting
Traditionally, the catastrophic forgetting (Kemker et al., 2018) refers to the phenomenon in deep learning where a neural network forgets previously learned knowledge when trained on a new task. To address this issue, researchers have proposed several methods, including: Synaptic Intelligence (SI) (Zenke et al., 2017) dynamically estimates the importance of each parameter in an online fashion, penalizing significant changes to parameters that are important for previously learned tasks during training on new tasks. This method adjusts the learning rate for parameters, ensuring that important parameters are not excessively modified. Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2016) grounded in a Bayesian perspective, estimates the importance of parameters by calculating the Fisher Information Matrix. During new task training, EWC introduces a regularization term that restricts the updates to important parameters, thereby preventing catastrophic forgetting. From a probabilistic viewpoint, EWC derives an importance matrix that quantifies the significance of network parameters for previous tasks. Learning without Forgetting (LwF) (Li & Hoiem, 2017) addresses catastrophic forgetting by incorporating outputs from previous tasks during training on a new task. This approach, based on knowledge distillation, transfers knowledge from old tasks to the new task, allowing the model to retain performance on the previous task when learning.
@@ 2.2 Catastrophic Forgetting in LLM and LoRA
With the rapid advancement of large language models (LLMs) (Touvron et al., 2023a, b), directly using pretrained models for domain-specific tasks has become prohibitively expensive. As a result, fine-tuning has become the preferred approach, typically divided into full parameter tuning and parameter-efficient fine-tuning (PEFT) methods, such as LoRA (Low-Rank Adaptation) (Hu et al., 2021; Wang et al., 2024). Full parameter fine-tuning (Lv et al., 2023) updates all model parameters to improve task adaptability but often causes catastrophic forgetting. PEFT methods like LoRA, by updating only a small subset of parameters through low-rank matrices, reduce computational costs and mitigate forgetting, though some still occurs.
To further reduce catastrophic forgetting, researchers have proposed combining EWC with LoRA in a method known as EWCLoRA (Xiang et al., 2024). This method leverages EWC to calculate the Fisher Information Matrix for parameter importance and uses low-rank matrices of LoRA to limit the scope of parameter updates. However, the calculation of the Fisher matrix introduces significant computational and memory overhead. Additionally, an interpolation-based LoRA (I-LoRA) method is introduced by Ren et al.. I-LoRA constructs a dual-memory experience replay framework, utilizing LoRA parameter interpolation to simulate the weight interpolation process. This approach strikes a balance between rapidly adapting to new tasks and retaining old knowledge. However, it requires maintaining an additional set of LoRA parameters throughout the process, which again increases the space cost.
@@ 3 Preliminary
LoRA is a lightweight and parameter-efficient fine-tuning method that introduces low-rank decomposition into the weight matrix $\theta$ of a pretrained model. Only the newly added low-rank matrices $B$ and $A$ are optimized, while the main weight $\theta_{0}$ remains frozen. The parameter at time $t$ during fine-tuning can be expressed as:
$\theta_{t}=\theta_{0}+\delta_{t},\quad\delta_{t}=B_{t}A_{t}$ (1)
where $\theta_{0}\in\mathbb{R}^{d\times d}$ are pretrained weights; $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times d}$ are the low-rank matrices with $r\ll d$.
@@
Figure 2: Adaptive constraint combining element-wise and layer-wise importance to preserve general capabilities from the $\nu$ task while learning domain-specific abilities for the $\mu$ task.
4 Adaptive Importance Finetuning
Inspired by Synaptic Intelligence (SI) (Zenke et al., 2017), we propose a method to constrain LLMs from making significant changes to their general capabilities during fine-tuning, thus addressing catastrophic forgetting. As shown in Figure 2, the main idea is to compute the importance of each parameter during the training of the initial general task (e.g. $\nu$) and constrain their updates when fine-tuning on subsequent tasks (e.g. $\mu$). Specifically, the importance scores measure how much each parameter contributes to reducing the loss in the $\nu$ task, and these scores are used to guide the fine-tuning process for the new $\mu$ task. This ensures that the critical parameters for $\nu$ task are modified to a lesser extent when learning $\mu$ task.
@@ 4.1 General Element-Wise Importance Recording
We define the contribution of parameter $i$ to the reduction of the loss function as $\omega_{i}$. The larger the value of $\omega_{i}$, the more important the parameter $i$ is for maintaining the performance of the task. The change in the loss function from time $t_{0}$ to time $t_{1}$ can be defined as the sum of the contributions of all parameters:
@@
In accordance with Equation 4 and Equation 6, it is deduced that the defined quantity $\omega_{k}$ corresponds precisely to the negative of the path integral of the gradient $g_{k}$.
@@
This indicates that we can represent $\omega_{i}$ using the product of $g(\theta_{i}(t))=\frac{\partial L}{\partial\theta_{i}}$ (subsequently simplified to $g_{i}(t)$) and $\theta_{i}^{\prime}(t)=\frac{\partial\theta_{i}}{\partial t}$ (Zenke et al., 2017).
@@
$A(t+1)=A(t)-\eta g^{A}(t)$ (9)
where $\eta$ is the learning rate, $g^{A}(t)$ and $g^{B}(t)$ are the gradients of the loss functions with respect to A and B. Based on Equation 8 and Equation 9, we derive the following expression:
@@
Based on Equation 12 and the definition of $\tilde{g}(t)$, we obtain the change in parameter value and hypothetical gradient at time t.
@@
In this way, we obtained the value of $\omega_{i}$ for the LoRA scenario.
@@
where $\Delta_{i}^{\nu}=\theta_{i}(t^{\nu})-\theta_{i}(t^{0})$ is whole change of the $i$-th parameter $\theta_{i}$ during task $\nu$, $\theta_{i}(t^{\nu})$ is the final parameter after fine-tuning on task $\nu$. In the context of LoRa fine-tuning, the $\Delta_{i}^{\nu}$ is defined as $B_{i}(t^{\nu})A_{i}(t^{\nu})$. This relationship stems from the fact that, at the initialization of LoRa at time 00, the B matrix is initialized to zero. The term in the denominator $(\Delta_{i}^{\nu})^{2}$ ensures that the regularization term carries the same units as the loss $L$. $\xi$ is a small positive constant to prevent division by zero.
This formulation assigns higher scores to parameters that have a significant impact on loss reduction while accounting for their magnitude to avoid bias toward large updates.
@@ 4.2 Element-Wise Regularization in Domain Tuning
After fine-tuning on the $\nu$ task, we extend the optimization objective to include both the task-specific loss and an SI-inspired regularization during $\mu$ finetuning. The task-specific loss $\mathcal{L}_{\text{task}}^{\mu}$ drives the adaptation to the $\mu$ task. To preserve knowledge from the $\nu$ task, the SI regularization term penalizes deviations from the important parameter values recorded in the $\nu$ task. The regularization term of the $l$-th layer is defined as:
$\mathcal{L}_{\text{reg},l}^{\nu}=\sum_{i\in p_{l}}\sum_{\nu<t<\mu}\Omega_{i}^{\nu}(\theta_{i}^{t}-\theta_{i}^{\nu})^{2}$ (16)
Here, $\Omega_{i}^{\nu}$ represents the importance of the $i$-th parameter in the $\nu$ task, and $\theta_{i}^{\nu}$ is the reference parameter after $\nu$ task fine-tuning. $p_{l}$ is the parameters of $l$-th layer. This term ensures that parameters with high importance scores remain close to their $\nu$ task values while allowing less important parameters more flexibility for adaptation.
During training, $\omega_{i}$ values are updated continuously, while the cumulative importance measures $\Omega_{i}^{\nu}$ is updated only at the end of task $\nu$. After updating $\Omega_{i}^{\nu}$, the $\omega_{i}$ values are reset to zero.
4.3 Layer-Wise Weighted Regularization

### v2

# How to Alleviate Catastrophic Forgetting in LLMs Finetuning? Hierarchical Layer-Wise and Element-Wise Regularization

@@ Abstract
Large Language Models (LLMs) exhibit strong general language capabilities. However, fine-tuning these models on domain-specific tasks often leads to catastrophic forgetting, where the model overwrites or loses essential knowledge acquired during pretraining. This phenomenon significantly limits the broader applicability of LLMs. To address this challenge, we propose a novel approach to compute the element-wise importance of model parameters crucial for preserving general knowledge during fine-tuning. Our method utilizes a dual-objective optimization strategy: (1) regularization loss based on element-wise parameter importance, which constrains the updates to parameters crucial for general knowledge; (2) cross-entropy loss to adapt to domain-specific tasks. Additionally, we introduce layer-wise coefficients to account for the varying contributions of different layers, dynamically balancing the dual-objective optimization. Extensive experiments on scientific, medical, and physical tasks using GPT-J and LLaMA-3 demonstrate that our approach mitigates catastrophic forgetting while enhancing model adaptability. Compared to previous methods, our solution is approximately 20 times faster and requires only 10%–15% of the storage, highlighting the practical efficiency. The code will be released.
@@ 1 Introduction
Large Language Models (LLMs) are pretrained on massive and diverse datasets, equipping them with remarkable general capabilities (Wang & Komatsuzaki, 2021; Touvron et al., 2023b; OpenAI, 2024). This pretraining process allows LLMs to serve as versatile tools for a wide range of natural language processing tasks. However, in domains such as medical and scientific fields, LLMs often struggle to perform effectively, necessitating fine-tuning domain-specific data. While fine-tuning could enhance the model task-specific performance, it also introduces a critical challenge: catastrophic forgetting (Kirkpatrick et al., 2016; Kemker et al., 2018; Shao & Feng, 2022; Ren et al., 2024). As shown in Figure 1, catastrophic forgetting refers to the phenomenon where a model, during the process of fine-tuning, loses or overwrites knowledge learned during pretraining. This issue poses a severe limitation on the broader applicability of LLMs, as it undermines their versatility and reusability across domains. The fixed data composition and format in the fine-tuning data may impair the general knowledge previously learned by the model. This results in a loss of logical reasoning abilities and related general knowledge, which affects the model performance on domain-specific tasks. On the other hand, it may also lead to a decline in the ability to answer general tasks, including questions it was previously capable of answering.
@@
Addressing catastrophic forgetting is therefore a crucial requirement for maximizing the utility of LLMs. A successful solution needs to achieve a delicate balance: retaining the essential general knowledge when learning new domain-specific expertise. This balance is critical when fine-tuning LLMs for specialized tasks, as both domain adaptation and generalizability are necessary for practical applications. EWCLoRA (Xiang et al., 2024) focuses on the issue of catastrophic forgetting in LLM fine-tuning and uses the Fisher matrix to measure the importance of parameters for general capabilities. However, it requires gradients computed with labels from the model distribution, necessitating an additional backpropagation pass for online computation. Thus, its computational cost is very high. For GPT-J-6B, calculating the Fisher matrix takes 22 hours on an A800 and requires 23GB of storage, and these requirements increase for larger LLMs. Besides, rsLoRA (Kalajdzievski, 2023) aims to stabilize learning by introducing a rank-stabilized scaling factor, but it does not effectively protect general capabilities as expected.
@@
To address catastrophic forgetting, we calculate parameter importance from two dimensions—element-wise and layer-wise—to constrain the updates of parameters crucial for general capabilities during fine-tuning. Firstly, our approach calculates the path integral during parameter updates as the element-wise parameter importance for regularization on the general capabilities of the LLM. This helps preserve parameters critical for general knowledge, minimizing significant modifications to it. Our method could avoid the computation and storage of the Fisher matrix, enabling faster and more storage-efficient computation of parameter importance. Specifically, we define domain $\nu$ as the general knowledge, representing the general capabilities of LLMs, and domain $\mu$ as the knowledge learned during fine-tuning for specific tasks. Our approach leverages a dual-objective optimization strategy that combines two losses: regularization loss, which reduces updates to parameters critical for domain $\nu$ to preserve general knowledge; and cross-entropy (CE) loss, which encourages alignment of domain $\mu$ parameters to enhance domain-specific learning. Through the constraint of a dual-objective loss, we aim to maintain general capabilities while performing domain-specific fine-tuning.
@@
Besides, we propose a layer-wise coefficient to adjust the weight of regularization loss. In LLMs, different layers contribute differently to generalization ability and domain-specific ability. The impact of each layer on the learning process is not uniform; some layers capture high-level abstract features, while others focus on lower-level details. Traditional approaches often treat the importance of each layer as equal, which overlooks the varying degrees of influence that different layers have on the model learning and generalization ability. Thus, we propose layer-wise coefficients to dynamically adjust the balance between task learning and the retention of general knowledge in each layer, allowing some layers to prioritize task learning, while others preserve general knowledge. We use the L2 norm of the computed element-wise importance of each layer weight to capture their contribution to both objectives.
@@
  • We introduce a framework that first records parameter importance on general data, and then applies regularization constraints during fine-tuning on domain-specific data to effectively address catastrophic forgetting in large language models (LLMs).
@@
  • Our method achieves state-of-the-art performance across multiple datasets using mainstream backbone LLMs. It significantly reduces computational time (20x faster) and storage (10$\sim$15%) for parameter importance estimation compared to prior methods.
@@ 2 Related Work
2.1 Continual Learning
Traditionally, continual learning (Wickramasinghe et al., 2024; Hadsell et al., 2020; Wickramasinghe et al., 2024; Vijayan & Sridhar, 2021) refers to developing learning algorithms to accumulate knowledge on non-stationary data. In general, continuous learning could be categorized into the following methods: Regularization-based methods. Synaptic Intelligence (SI) (Zenke et al., 2017) dynamically estimates the importance of each parameter in an online fashion, penalizing significant changes to parameters that are important for previously learned tasks during training on new tasks. This method adjusts the learning rate for parameters, ensuring that important parameters are not excessively modified. Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2016) grounded in a Bayesian perspective, estimates the importance of parameters by calculating the Fisher Information Matrix. During new task training, EWC introduces a regularization term that restricts the updates to important parameters, thereby preventing catastrophic forgetting. From a probabilistic viewpoint, EWC derives an importance matrix that quantifies the significance of network parameters for previous tasks. Architecture-based methods. Researches (Wu et al., 2024; Wang et al., 2023, 2022; Chen et al., 2022) learn new tasks by adapting the structure of existing models. For instance, Wang et al. inserts trainable task-specific prompts to the input layer to expand the domain ability. Replay-based methods. Researchers (Jin et al., 2022; Liu et al., 2021; Qin et al., 2022; Bai et al., 2022) retain a subset of previously encountered data, which are reintegrated into the training process of the new tasks. Distillation-based methods. Researches (Li & Hoiem, 2017; Cao et al., 2021; Shao & Feng, 2022; Gu et al., 2022; Qin & Joty, 2022) learn new tasks under the guidance of a teacher model. For instance, Learning without Forgetting (LwF) (Li & Hoiem, 2017) transfers knowledge from old tasks to new tasks, allowing the model to retain performance on the previous task while learning new ones.
@@ 2.2 Catastrophic Forgetting in LLM and LoRA
With the rapid advancement of large language models (LLMs) (Touvron et al., 2023a, b), directly using pretrained models for domain-specific tasks has become prohibitively expensive. As a result, fine-tuning has become the preferred approach, typically divided into full parameter tuning and parameter-efficient fine-tuning (PEFT) methods, such as LoRA (Low-Rank Adaptation) (Hu et al., 2021; Wang et al., 2024). Full parameter fine-tuning (Lv et al., 2023) updates all model parameters to improve task adaptability but often causes catastrophic forgetting. PEFT methods like LoRA, by updating only a small subset of parameters through low-rank matrices, reduce computational costs and mitigate forgetting, though some still occur.
To further reduce catastrophic forgetting, researchers have proposed combining EWC with LoRA in a method known as EWCLoRA (Xiang et al., 2024). This method leverages EWC to calculate the Fisher Information Matrix for parameter importance and uses low-rank matrices of LoRA to limit the scope of parameter updates. However, the calculation of the Fisher matrix introduces significant computational and memory overhead. Additionally, an interpolation-based LoRA (I-LoRA) method is introduced by Ren et al.. I-LoRA constructs a dual-memory experience replay framework, utilizing LoRA parameter interpolation to simulate the weight interpolation process. However, it requires maintaining an additional set of LoRA parameters throughout the process, increasing space cost.
@@ 3 Preliminary
LoRA is a lightweight and parameter-efficient fine-tuning method that introduces low-rank decomposition into the weight matrix $\theta$ of a pretrained model. Only the newly added low-rank matrices $B$ and $A$ are optimized, while the main weight $\theta_{0}$ remains frozen. The parameter at time $t$ during fine-tuning can be expressed as $\theta_{t}=\theta_{0}+\delta_{t};\delta_{t}=B_{t}A_{t}$, where $\theta_{0}\in\mathbb{R}^{d\times d}$ are pretrained weights; $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times d}$ are the low-rank matrices with $r\ll d$.
@@
Figure 2: Adaptive constraint combining element-wise and layer-wise importance to preserve general capabilities from the $\nu$ task while learning domain-specific abilities for the $\mu$ task. RECORD means the general importance recording in Section 4.1. REG means the regularization in Section 4.2 and Section 4.3.
4 Hierarchical Importance Regularization
Inspired by Synaptic Intelligence (SI) (Zenke et al., 2017), we propose a framework to constrain LLMs from making significant changes to their general capabilities during fine-tuning, thus addressing catastrophic forgetting in LoRA tuning. As shown in Figure 2, the framework is to compute the importance of each parameter during the training of the initial general task (e.g. $\nu$) and constrain their updates when fine-tuning on subsequent task (e.g. $\mu$). Specifically, the importance scores measure how much each parameter contributes to reducing the loss in the $\nu$ task, and these scores are used to guide the fine-tuning process for the new $\mu$ task. This ensures that the critical parameters for $\nu$ task are modified to a lesser extent when learning $\mu$ task.
@@ 4.1 General Element-Wise Importance Recording
We define the contribution of parameter $i$ to the reduction of the loss function as $\omega_{i}$. The larger the value of $\omega_{i}$, the more important the parameter $i$ is for maintaining the performance of the task $\nu$. The change in the loss function from time $t_{0}$ to time $t_{1}$ can be defined as the sum of the contributions of all parameters:
@@
In accordance with Equation 3 and Equation 5, it is deduced that the defined quantity $\omega_{i}$ corresponds precisely to the negative of the path integral of the gradient $g_{i}$.
@@
This indicates that we can represent $\omega_{i}$ using the product of $g(\theta_{i}(t))=\frac{\partial L}{\partial\theta_{i}}$ and $\theta_{i}^{\prime}(t)=\frac{\partial\theta_{i}}{\partial t}$ (Zenke et al., 2017).
@@
$A(t+1)=A(t)-\eta g^{A}(t)$
where $\eta$ is the learning rate, $g^{A}(t)$ and $g^{B}(t)$ are the gradients of the loss functions with respect to A and B. Based on Equation 7, we derive the following expression:
@@
Based on Equation 10 and the definition of $\tilde{g}(t)$, we obtain the parameter change and hypothetical gradient at time t.
@@
In this way, we obtain the value of $\omega_{i}$ for the LoRA scenario.
@@
where $\Delta_{i}^{\nu}=\theta_{i}(t^{\nu})-\theta_{i}(t^{0})$ is whole change of the $i$-th parameter $\theta_{i}$ during task $\nu$, $\theta_{i}(t^{\nu})$ is the final parameter after fine-tuning on task $\nu$. In the context of LoRA fine-tuning, the $\Delta_{i}^{\nu}$ is defined as $(B(t^{\nu})A(t^{\nu}))_{i}$. This relationship stems from the fact that, at the initialization of LoRA at time 00, the B matrix is set to zero. The term in the denominator $(\Delta_{i}^{\nu})^{2}$ ensures that the regularization term carries the same units as the loss $L$. $\xi$ is a small positive constant to prevent division by zero. This formulation assigns higher scores to parameters that have a significant impact on loss reduction while accounting for their magnitude to avoid bias toward large updates.
@@ 4.2 Element-Wise Regularization in Domain Tuning
After fine-tuning the $\nu$ task, we extend the optimization objective to include both task-specific and regularization losses during $\mu$ finetuning. The task-specific loss $\mathcal{L}_{\text{task}}^{\mu}$ drives the adaptation to the $\mu$ task. To preserve knowledge from the $\nu$ task, the regularization loss penalizes deviations from the important parameter values recorded in the $\nu$ task. The regularization loss $\mathcal{L}_{\text{reg},l}^{\nu}$ of the $l$-th layer is defined as:
$\mathcal{L}_{\text{reg},l}^{\nu}=\sum_{i}^{n}\sum_{\nu<t<\mu}\Omega_{i}^{\nu}(\theta_{i}^{t}-\theta_{i}^{\nu})^{2}$ (14)
where $\Omega_{i}^{\nu}$ represents the importance of the $i$-th parameter in the $\nu$ task, and $\theta_{i}^{\nu}$ is the reference parameter after $\nu$ task fine-tuning. This loss ensures that parameters with high importance scores remain close to their $\nu$ task values while allowing less important parameters more flexibility for adaptation. During training, $\omega_{i}$ values are updated continuously, while the cumulative importance $\Omega_{i}^{\nu}$ is updated only at the end of task $\nu$. After updating $\Omega_{i}^{\nu}$, the $\omega_{i}$ is reset to zero.
4.3 Layer-Wise Coefficient Regularization